OpenStudy (anonymous):

Find an integer which is the limit of (1-cosx)/x as x goes to 0

OpenStudy (yuki):

you can use the sandwitch thm

OpenStudy (amistre64):

1/x - cosx/x

OpenStudy (yuki):

or you can do L'Hospitals rule

OpenStudy (amistre64):

hopitals rule was my thought :)

OpenStudy (yuki):

if you take derivative of both top and bottom you'll get sin(x)

OpenStudy (yuki):

so as x-->0, sin(x)-->0 so the answer is 0

OpenStudy (amistre64):

we can always do the long version :) 1-cos(x+h) - 1 + cos(x) --------------------- maybe? h

OpenStudy (anonymous):

thank you all. it's 0 as yuki said hehe

OpenStudy (amistre64):

its always been 0 lol

OpenStudy (yuki):

is you use the sandwich thm \[-1 \le \cos(x) \le 1 \] \[1 \ge -\cos(x) \ge -1\] \[2 \ge 1-\cos(x) \ge 0\] \[{2 \over x } \ge {1-\cos(x) \over x} \ge 0\]

OpenStudy (yuki):

now if you take the limit on both sides \[\lim_{x \rightarrow 0} {2 \over x} \ge \lim_{x \rightarrow 0}{1-\cos(x) \over x} \ge \lim_{x \rightarrow 0}{0}\] \[0 \ge \lim_{x \rightarrow 0}{1-\cos(x) \over x} \ge 0\] so the limit is 0